OpenStudy (anonymous):

Los Angeles is about 385 miles from San Fransico. How far spat would the cities be on a map with a scale of 1in.=mi? Round to the nearest hundredth

OpenStudy (mathstudent55):

What is the scale? Is there a number missing next to mi?

OpenStudy (anonymous):

Oh it's supposed to be 1 in. = 25 mi

OpenStudy (mathstudent55):

Ok. You can do scale problems with a proportion. If 1 in. equals 25 mi, then how many inches equals 385 mi? A proportion is two equal ratios. 1 in. is to 25 miles as x inches is to 385 miles. \(\dfrac{1~in.}{25~mi} = \dfrac{x}{385~mi} \) Cross multiply: \(25x = 385\times 1\) \(25x = 385\) Divide both sides by 25: \(x = \dfrac{385}{25} \) \(x = 15.4 ~in.\)